# Second Law of Entanglement Manipulation with Entanglement Battery

### v1

@@ Abstract
We emphasize that Theorem 3 is fully general, i.e. it holds for generic states, as opposed to most results in single-shot thermodynamics that hold only for energy-incoherent states. Furthermore, we can even allow the system and battery to get correlated, similar to the entanglement case. Because free energy is a superadditive measure, it will still govern the allowed transformations. The proof of Theorem 3 relies on the key insight that thermal operations allow for the swapping of system and battery states. To reiterate, while swapping only the state between two systems is only allowed if they have the same Hamiltonian, swapping both the state and the Hamiltonian is always allowed since it amounts to a relabelling of the systems.
Note that as we take the limit of the temperature $T\to\infty$, the free energy of any state diverges except for the Gibbs state. The condition in Theorem 3 suggests that in this limit, the set of states can be divided into two classes: the Gibbs state and everything else. However, this is not the case; if we allow the battery to have infinite free energy, then we can subtract any finite amount of free energy from the battery and satisfy the resource non-decreasing condition on the battery. This would allow us to create any state, even starting from the Gibbs state. Thus, in order to obtain meaningful conditions for state manipulations, the resource measure of the battery should be finite.
@@
which hold for any additive monotone $f$. However, when we allow correlations, then Eq. (8) is equivalent to the local monotonicity property studied in Ref. [45]. There, it was shown that standard free energy $F$ is essentially the only function that is locally monotonic and continuous, up to additive and multiplicative constants. This gives a formulation of the thermodynamic second law for coherent quantum systems.
It is worth noting that the state transformation conditions in Theorem 3 are identical to that of Gibbs-preserving operations with correlated catalysis [33]. Thus, our theorem provides an operational interpretation of catalytic Gibbs-preserving operations as thermal operations augmented with a free energy battery. In our setting, the battery can act as a reservoir of coherence that can be freely used. At the same time, we can prevent pumping free energy into the system by constraining the free energy of the battery. It would be interesting to investigate whether we can augment thermal operations in a similar way to reproduce the non-catalytic version of Gibbs-preserving operations.
Conclusions
In this article, we have explored the concept of entanglement batteries and their impact on the manipulation of entangled systems. Our findings illuminate the path toward achieving reversible entanglement manipulations across all quantum states, thereby addressing a long-standing challenge in quantum information science.
@@ Conclusions
This work also opens several avenues for future research, presenting intriguing questions pivotal for a deeper understanding of entangled systems. A particularly compelling area for further investigation involves identifying all entanglement quantifiers that yield conversion rates in the form of $E(\rho)/E(\sigma)$. Although it was previously speculated [4] that the regularized relative entropy of entanglement might uniquely possess this characteristic, our findings hint at the possibility that other entanglement quantifiers may also be suitable for this task.
@@
Methods
LOCC with entanglement battery
@@ Methods
In the setting defined above, the battery is not correlated with the main system at the end of the procedure. In the Supplemental Material, we also discuss the more general setting where correlations between the main system and the battery are taken into account.
Proof of Theorems 1 and 2
In the following, we assume that $E$ is an additive and asymptotically continuous measure [26], i.e.,
@@
To prove Theorem 1, let us first assume that $\rho^{AB}$ can be converted into $\sigma^{AB}$ via LOCC with entanglement battery. Then, there is an LOCC protocol $\Lambda$ and states $\tau^{A^{\prime}B^{\prime}}$ and $\tilde{\tau}^{A^{\prime}B^{\prime}}$ such that Eqs. (9) and  (10) are fulfilled. Using the additivity of $E$, finiteness and its monotonicity under LOCC, we obtain
@@
A conversion $\rho\rightarrow\sigma$ can be achieved in this setting by choosing
@@
Due to Eq. (15) we have $E(\tilde{\tau})\geq E(\tau)$ as required, and thus this protocol achieves the transformation $\rho^{AB}\rightarrow\sigma^{AB}$. This completes the proof of Theorem 1.
We will now prove Theorem 2, showing that for any entanglement measure which is additive and asymptotically continuous, the asymptotic transformation rates take the form
@@
$\frac{m}{n}\leq\frac{E(\rho)+\frac{f(\varepsilon)}{n}}{E(\sigma)-\varepsilon K\log_{2}d}.$ (23)
Assuming that
$0<\varepsilon<\frac{E(\sigma)}{K\log_{2}d}$ (24)
and using Eqs. (11), we see that any feasible rate $r$ must fulfill
@@
Proof of Theorem 3
@@
Zero error conversion
@@
Entanglement measures which are additive and asymptotically continuous
As mentioned in the main text, an example for an entanglement measure which is additive and asymptotically continuous is the squashed entanglement given in Eq. (5) with the quantum conditional mutual information of a state $\rho^{ABE}$ defined as
@@
Other entanglement measures
@@
where $\mathcal{S}$ is the set of separable states, and $S(\rho||\sigma)=\mathrm{Tr}(\rho\log_{2}\rho)-\mathrm{Tr}(\rho\log_{2}\sigma)$ is the quantum relative entropy. Since the relative entropy of entanglement is not additive [51], our Theorems do not directly apply in this case. Nevertheless, as we show in the Supplemental Material, quantifying the amount of entanglement in the battery by the relative entropy of entanglement leads to a theory with bounded distillation rates, in particular the asymptotic rate for converting a state $\rho$ into |$\ket{\psi^{-}}$ is upper bounded by $E_{\mathrm{r}}(\rho)$. Moreover, we show that $E_{\mathrm{r}}^{\infty}(\rho)/E_{\mathrm{r}}^{\infty}(\sigma)$ is a feasible transformation rate for any two entangled states $\rho$ and $\sigma$, where $E_{\mathrm{r}}^{\infty}$ is the regularized relative entropy of entanglement defined in Eq. (6). The same rate is also feasible if the amount of entanglement in the battery is quantified via $E_{\mathrm{r}}^{\infty}$. In this setting, we also obtain a theory with bounded distillation rates. We note that for some states $\rho$ and $\sigma$ the asymptotic transformation rate might exceed $E_{\mathrm{r}}^{\infty}(\rho)/E_{\mathrm{r}}^{\infty}(\sigma)$. We refer to the Supplemental Material for a more detailed discussion.
Another frequently used entanglement measure in the literature is the logarithmic negativity [28, 29], defined as
@@
where $T_{A}$ denotes partial transposition. Although this measure is additive, it does not satisfy asymptotic continuity [7]. As discussed above in the Methods section, the additivity of $E_{\mathrm{n}}$ implies a reversible framework for zero error transformations. Additionally, in the Supplemental Material, we demonstrate that the entanglement distillation rate is bounded even when allowing for an error margin in the transformation. Specifically, the asymptotic rate for the conversion $\rho\rightarrow\ket{\psi^{-}}$ is given by $E_{\mathrm{n}}(\rho)$. We further note that logarithmic negativity vanishes if and only if the state has positive partial transpose (PPT), which means that this framework allows for the creation of PPT states from separable states.
@@
First, recall that we can convert any state $\rho$ into singlets at rate $E_{\mathrm{n}}(\rho)$ in this setup. As discussed above in the Methods section, this conversion can be achieved with zero error. Moreover, by using LOCC it is possible to convert singlets approximately into the state $\rho$ at rate $1/E_{\mathrm{c}}(\rho)$, where $E_{\mathrm{c}}$ is the entanglement cost [52]. Thus, performing these operations in sequence will convert $\rho^{\otimes n}$ into a state which is close to $nE_{\mathrm{n}}(\rho)/E_{\mathrm{c}}(\rho)$ copies of $\rho$, and the error can be made arbitrarily small in the limit of large $n$. Self-dilution of $\rho$ occurs whenever $E_{\mathrm{n}}(\rho)/E_{\mathrm{c}}(\rho)>1$.
@@
This phenomenon is not unique to logarithmic negativity, or even entanglement theory. By analogous arguments, we can show that self-dilution occurs in thermodynamics for incoherent states if we use $F_{\textrm{max}}(\rho)=\inf\Bqty{\log\lambda\,|\,\rho\leq\lambda\gamma}$ to quantify the resources in the battery. Similarly to logarithmic negativity, $F_{\textrm{max}}(\rho)$ is not asymptotically continuous and has the operational interpretation as the exact cost of preparing $\rho$ from many copies of |$\outerproduct{1}{1}$ | [38]. We refer to the Supplementary Material for more details.
@@
Acknowledgements
Supplemental Material
Connection to catalysis
$\Lambda\pqty{\rho^{S}\otimes\tau^{C}}$ ) $=\sigma^{S}\otimes\tau^{C}.$ (37)
Relative entropy of entanglement
@@ Acknowledgements
Logarithmic negativity
We will now quantify entanglement using logarithmic negativity [28, 29]
@@
More details can also be found in the Methods section.
@@
Correlations between main system and battery
We will now consider a more general setting where the battery might become correlated with the main system. We say that a state $\rho$ can be transformed into another state $\sigma$ in this setting if for any $\varepsilon>0$ there exists an LOCC protocol $\Lambda$ and a state of the battery $\tau$ such that
@@
Note that asymptotic continuity is not required in the following. Examples of entanglement measures fulfilling these properties are the squashed entanglement [27] and the conditional entanglement of mutual information [49, 50] (see main text and Methods section for the definition of these measures). To see that conditional entanglement of mutual information fulfills superadditivity (property 2), consider an extension $\tau^{AA^{\prime}A^{\prime\prime}BB^{\prime}B^{\prime\prime}}$ of the state $\rho^{AA^{\prime}BB^{\prime}}$. It holds that
@@
Self-dilution in thermodynamics
@@
Now, consider the following process: we start with $\rho^{\otimes n}$ and we distill |$\outerproduct{1}{1}^{\otimes nr}$ with the help of a battery, where $r=F_{\textrm{max}}(\rho)/F_{\textrm{max}}(\outerproduct{1}{1})$. Since we are doing this with a battery, we know that this can be done without error. Then, we perform a dilution procedure |$\outerproduct{1}{1}^{\otimes nr}\to\rho^{\otimes nrr^{\prime}}$ without battery, with $r^{\prime}=F(\outerproduct{1}{1})/F(\rho)$ [65]. Then, we have

### v2

@@ Abstract
We emphasize that Theorem 3 holds for generic states, as opposed to most results in single-shot thermodynamics that hold only for energy-incoherent states. It is worth noting that the state transformation conditions in Theorem 3 are identical to that of Gibbs-preserving operations with correlated catalysis [28]. Thus, our theorem provides an operational interpretation of catalytic Gibbs-preserving operations as thermal operations augmented with a free energy battery.
@@
which hold for any additive monotone $f$. However, when we allow correlations, then Eq. (8) is equivalent to the local monotonicity property studied in Ref. [42]. There, it was shown that the standard free energy $F$ is essentially the only function that is locally monotonic and continuous, up to additive and multiplicative constants. Combined with our results, this gives a formulation of the unique thermodynamic second law for coherent quantum systems.
Conclusions. In this article, we have explored the concept of entanglement batteries and their impact on the manipulation of entangled systems. Our findings illuminate the path toward achieving reversible entanglement manipulations across all quantum states, thereby addressing a long-standing challenge in quantum information science.
@@
This work opens several avenues for future research, presenting questions pivotal for a deeper understanding of entangled systems. A particularly compelling area for further investigation involves identifying all entanglement quantifiers that yield conversion rates in the form of $E(\rho)/E(\sigma)$. It was previously speculated [2] that the regularized relative entropy of entanglement might uniquely possess this characteristic, but our findings hint at the possibility that other entanglement quantifiers may also be suitable for this task.
@@
Note added. After the completion of our work, two proofs of the generalized quantum Stein’s lemma were posted on arXiv [44, 45], establishing the reversibility of entanglement theory under asymptotically non-entangling operations. These results are complementary to our work, and show that reversibility of entanglement can be obtained in multiple frameworks.
End Matters
.1 LOCC with entanglement battery
@@
In the setting defined above, the battery is not correlated with the main system at the end of the procedure. In the Supplemental Material, we discuss the more general setting where correlations between the main system and the battery are taken into account.
.2 Proof of Theorems 1 and 2 of the main text
We begin with the assumption that $E$ is an additive and asymptotically continuous measure [23], i.e.,
@@
To prove Theorem 1 of the main text, let us first assume that $\rho^{AB}$ can be converted into $\sigma^{AB}$ via LOCC with entanglement battery. Then, there is an LOCC protocol $\Lambda$ and states $\tau^{A^{\prime}B^{\prime}}$ and $\tilde{\tau}^{A^{\prime}B^{\prime}}$ such that Eqs. (9) and  (10) are fulfilled. Using the additivity of $E$, finiteness and its monotonicity under LOCC, we obtain
@@
A conversion $\rho\rightarrow\sigma$ can be achieved by choosing
@@
Due to Eq. (15) we have $E(\tilde{\tau})\geq E(\tau)$ as required, and thus this protocol achieves the transformation $\rho^{AB}\rightarrow\sigma^{AB}$. This completes the proof of Theorem 1 of the main text.
We will now prove Theorem 2 of the main text, showing that for any entanglement measure which is additive and asymptotically continuous, the asymptotic transformation rates take the form
@@
$\frac{m}{n}\leq\frac{E(\rho)+\frac{f(\varepsilon)}{n}}{E(\sigma)-\varepsilon K\log_{2}d},$ (23)
when
$0<\varepsilon<\frac{E(\sigma)}{K\log_{2}d}.$ (24)
Finally, using Eqs. (11), we see that any feasible rate $r$ must fulfill
@@
Supplemental Material
.3 Multiple measures on the battery
Let us explore what happens when we impose two non-equivalent [48] entanglement measures on the battery. By non-equivalent, we mean that we can find two states $\rho,\sigma$ such that $E_{1}(\rho)>E_{1}(\sigma)$, but $E_{2}(\rho)<E_{2}(\sigma)$. By following the arguments in Theorem 2, we can show that the transformation rates are bounded as follows
$R(\rho\to\sigma)$ $\leq\min\Bqty{\frac{E_{1}(\rho)}{E_{1}(\sigma)},\frac{E_{2}(\rho)}{E_{2}(\sigma)}}=\frac{E_{2}(\rho)}{E_{2}(\sigma)}<1,$ (28)
$R(\sigma\to\rho)$ $\leq\min\Bqty{\frac{E_{1}(\sigma)}{E_{1}(\rho)},\frac{E_{2}(\sigma)}{E_{2}(\rho)}}=\frac{E_{1}(\sigma)}{E_{1}(\rho)}<1,$ (29)
so that $R(\rho\to\sigma)R(\sigma\to\rho)<1$. This shows that the battery-assisted transformations only becomes reversible when we impose the non-decreasing condition on only one entanglement measure.
.4 Proof of Theorem 3 of the main text
@@
.5 Zero error conversion
@@
.6 Entanglement measures which are additive and asymptotically continuous
As mentioned in the main text, an example for an entanglement measure which is additive and asymptotically continuous is the squashed entanglement given in Eq. (5) of the main text, with the quantum conditional mutual information of a state $\rho^{ABE}$ defined as
@@
.7 Other entanglement measures
@@
where $\mathcal{S}$ is the set of separable states, and $S(\rho||\sigma)=\mathrm{Tr}(\rho\log_{2}\rho)-\mathrm{Tr}(\rho\log_{2}\sigma)$ is the quantum relative entropy. Since the relative entropy of entanglement is not additive [51], our Theorems do not directly apply in this case. Nevertheless, as we show below, quantifying the amount of entanglement in the battery by the relative entropy of entanglement leads to a theory with bounded distillation rates, in particular the asymptotic rate for converting a state $\rho$ into |$\ket{\psi^{-}}$ is upper bounded by $E_{\mathrm{r}}(\rho)$. Moreover, we show that $E_{\mathrm{r}}^{\infty}(\rho)/E_{\mathrm{r}}^{\infty}(\sigma)$ is a feasible transformation rate for any two entangled states $\rho$ and $\sigma$, where $E_{\mathrm{r}}^{\infty}$ is the regularized relative entropy of entanglement defined in Eq. (6) of the main text. The same rate is also feasible if the amount of entanglement in the battery is quantified via $E_{\mathrm{r}}^{\infty}$. In this setting, we also obtain a theory with bounded distillation rates. We note that for some states $\rho$ and $\sigma$ the asymptotic transformation rate might exceed $E_{\mathrm{r}}^{\infty}(\rho)/E_{\mathrm{r}}^{\infty}(\sigma)$.
Another frequently used entanglement measure in the literature is the logarithmic negativity [52, 53, 54], defined as
@@
where $T_{A}$ denotes partial transposition. Although this measure is additive, it does not satisfy asymptotic continuity [5]. As discussed above in the Supplemental Material, the additivity of $E_{\mathrm{n}}$ implies a reversible framework for zero error transformations. Below, we demonstrate that the entanglement distillation rate is bounded even when allowing for an error margin in the transformation. Specifically, the asymptotic rate for the conversion $\rho\rightarrow\ket{\psi^{-}}$ is given by $E_{\mathrm{n}}(\rho)$. We further note that logarithmic negativity vanishes if and only if the state has positive partial transpose (PPT), which means that this framework allows for the creation of PPT states from separable states.
@@
First, recall that we can convert any state $\rho$ into singlets at rate $E_{\mathrm{n}}(\rho)$ in this setup. As discussed above in the Supplemental Material, this conversion can be achieved with zero error. Moreover, by using LOCC it is possible to convert singlets approximately into the state $\rho$ at rate $1/E_{\mathrm{c}}(\rho)$, where $E_{\mathrm{c}}$ is the entanglement cost [55]. Thus, performing these operations in sequence will convert $\rho^{\otimes n}$ into a state which is close to $nE_{\mathrm{n}}(\rho)/E_{\mathrm{c}}(\rho)$ copies of $\rho$, and the error can be made arbitrarily small in the limit of large $n$. Self-dilution of $\rho$ occurs whenever $E_{\mathrm{n}}(\rho)/E_{\mathrm{c}}(\rho)>1$.
@@
This phenomenon is not unique to logarithmic negativity, or even entanglement theory. By analogous arguments, we can show that self-dilution occurs in thermodynamics for incoherent states if we use $F_{\textrm{max}}(\rho)=\inf\Bqty{\log\lambda\,|\,\rho\leq\lambda\gamma}$ to quantify the resources in the battery. Similarly to logarithmic negativity, $F_{\textrm{max}}(\rho)$ is not asymptotically continuous and has the operational interpretation as the exact cost of preparing $\rho$ from many copies of |$\outerproduct{1}{1}$ | [33]. This will be discussed in more detail below in the Supplemental Material
@@
.7.1 Relative entropy of entanglement
@@
.7.2 Logarithmic negativity
We will now quantify entanglement using logarithmic negativity [52, 53, 54]
@@
.8 Connection to catalysis
$\Lambda\pqty{\rho^{S}\otimes\tau^{C}}$ ) $=\sigma^{S}\otimes\tau^{C}.$ (86)
There is another connection to catalysis. Let us examine what happens to the battery state in a transformation cycle $\rho\to\sigma\to\rho$ when we perform the swapping protocol. Since $R(\rho\to\sigma)R(\sigma\to\rho)=1$, for any $\delta>0$, there is $m,n,s$ s.t. $\rho^{\otimes m}\to\sigma^{\otimes n}\to\rho^{\otimes s}$ and $s/m\geq 1-\delta$. With the swapping protocol, initially we prepare the battery in the state $\tau=\sigma^{\otimes n}\otimes\rho^{\otimes s}$. After the first stage of the transformation, the battery is in the state $\tau^{\prime}=\rho^{\otimes m}\otimes\rho^{\otimes s}$ and after the second stage, the state is $\tau^{\prime\prime}=\sigma^{\otimes n}\otimes\rho^{\otimes m}$. Notice that although the state of the battery after the first stage of the cycle $\tau^{\prime}$ can be far from its initial state $\tau$, the final state of the battery $\tau^{\prime\prime}$ is near its initial state $\tau$. This shows that the battery behaves almost catalytically in the whole cycle, although it might have large fluctuations in each individual stage.
.9 Correlations between main system and battery
We will now consider a more general setting where the battery might become correlated with the main system. We say that a state $\rho$ can be transformed into another state $\sigma$ in this setting if for any $\varepsilon>0$, there exists an LOCC protocol $\Lambda$ and a state of the battery $\tau$ such that
@@
Note that asymptotic continuity is not required in the following. Examples of entanglement measures fulfilling these properties are the squashed entanglement [24] and the conditional entanglement of mutual information [49, 50]. To see that conditional entanglement of mutual information fulfills superadditivity (property 2), consider an extension $\tau^{AA^{\prime}A^{\prime\prime}BB^{\prime}B^{\prime\prime}}$ of the state $\rho^{AA^{\prime}BB^{\prime}}$. It holds that
@@
.10 Self-dilution in thermodynamics
@@
Now, consider the following process: Starting with $\rho^{\otimes n}$, we distill |$\outerproduct{1}{1}^{\otimes nr}$ with the help of a battery, where $r=F_{\textrm{max}}(\rho)/F_{\textrm{max}}(\outerproduct{1}{1})$. With a battery, this transformation can be achieved with zero error. Then, we perform a dilution procedure |$\outerproduct{1}{1}^{\otimes nr}\to\rho^{\otimes nrr^{\prime}}$ without battery, with $r^{\prime}=F(\outerproduct{1}{1})/F(\rho)$ [69]. Then, we have